[Brahms-l] Shift report 20050605 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 05 2005 - 07:58:24 EDT
Shifter leader was rk
Shifters were 


1. Got beam all night but for 1.5 hour between stores (from 4.25 to 5.55).
2. Runs:
FS@2.3@861 B, MRS@52.5@4kG A:
15210 - 15217 -> 991035 tr2 ev., 478329 tr3 ev.
FS@2.3@861 A, MRS@40  @4kG A:
15220 - 15227 -> 281825 tr2 ev., 250676 tr3 ev.
15218,15219 - pedestals, 15228 running.
3. Still have no trigger 6, "scaledown calculator" is not working properly, no other problems observed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 5 07:58:37 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 05 2005 - 07:58:50 EDT